OpenStudy (anonymous):

Find x-y if 4x-3y-sqroot(4x-3y)= 5 and 7x-2y+sqroot(7x-2y)=1

OpenStudy (anonymous):

Let sqroot(4x-3y) = a a^2 - a = 5 Let sqroot(7x-2y) = b b^2 + b = 1

OpenStudy (anonymous):

\[b = 1/2 (-1\pm \sqrt{5})\]\[a = 1/2 (1\pm \sqrt{21})\]

OpenStudy (anonymous):

Doesnt that make x - y to be evaluated more complicated?

OpenStudy (anonymous):

But either way you can still get the same and correct answer